OpenStudy (anonymous):

Given a polynomial f(x), if (x − 2) is a factor, what else must be true?

zepdrix (zepdrix):

What else must be true? Hmm lots of stuff... this question is strangely broad. x=2 is a zero of the function. f(x)/(x-2) is a polynomial.

zepdrix (zepdrix):

Do you have some multiple choice options? :)

OpenStudy (anonymous):

f(0) = 2 f(0) = −2 f(−2) = 0 f(2) = 0

OpenStudy (anonymous):

I was thinking the first one

zepdrix (zepdrix):

So like I mentioned, x=2 is a zero of the function. That means when you replace x with 2, the function output is zero. A? I think you've got that backwards. In A, they've replaced x with 0 and the output is 2. f(x) -> f(0)

OpenStudy (anonymous):

Oh okay, my bad. So it's the last one. You substitute x for 0

zepdrix (zepdrix):

You substitute 2 for x... yes the last one :P

OpenStudy (anonymous):

Ugh yeah, that's I meant lol. I was still thinking of A. Thank you so much :)

OpenStudy (anonymous):

Do you mind helping me with another one?

OpenStudy (anonymous):

Which polynomial identity will prove that 49 = (2 + 5)2? Difference of Squares Difference of Cubes Sum of Cubes Square of a Binomial I think it's difference of squares

zepdrix (zepdrix):

Recall that in math, the word `difference` means `subtraction`. So that refers to some formula involving subtraction. No subtraction in our problem though, ya?

OpenStudy (anonymous):

Oh yeah. So it can't be A or B and this problem isn't a binomial so it's C?

zepdrix (zepdrix):

Well are you familiar with these words they're using? What is a cube? how large? How large is a square?

OpenStudy (anonymous):

A cube is like a 3d square, it can be any size but all sides need to be the same.

OpenStudy (anonymous):

Wait... if it's (2+5)^2 and you were to write it out; (2+5) * (2+5) that would be a binomial.

zepdrix (zepdrix):

3d is the key there. So in multiplication it's something being multiplied by itself 3 times. But we don't have that in our problem, ya? (2+5)^2 =/= (2+5)(2+5)(2+5) We don't have a cube.

zepdrix (zepdrix):

So square ya :) and yes a binomial is just two things being added or subtracted

OpenStudy (anonymous):

I knew it wasn't a cube because it was being squared, not cubed. Thank you again! :)

OpenStudy (anonymous):

Can you help me with this last one? This one I don't really know how to solve.

OpenStudy (anonymous):

Using a directrix of y = 5 and a focus of (4, 1), what quadratic function is created?

OpenStudy (anonymous):

Answer choices: f(x) = one fourth (x − 4)2 − 3 f(x) = one eighth (x + 4)2 − 3 f(x) = −one eighth (x − 4)2 + 3 f(x) = −one fourth (x + 4)2 − 3